OpenStudy (kaylala):

trigonometry: Identify if the underlined word/s is true or false If false change it to the appropriate one If the arc length is 30π_feet with central angle equal to (3/4)revolution. The radius is equal to 20feet. 20 is the underlined word

OpenStudy (kaylala):

here's a picture of the question: see item # 7

OpenStudy (***[isuru]***):

use .. \[S = r \ \theta\]

OpenStudy (***[isuru]***):

where S is the arc length r is the radius theta is the angle measure in radiant

OpenStudy (***[isuru]***):

angle of 3/4 revolution = 270 degrees = 3pi /2

OpenStudy (***[isuru]***):

and the length of the arc is 30 pi feet

OpenStudy (***[isuru]***):

\[S = r \ \theta\]\[30 \pi = r \times (\frac{ 3 \pi }{ 2 })\]

OpenStudy (***[isuru]***):

now make r the subject and c if the value u get for r is the same as in the question ..

OpenStudy (***[isuru]***):

got it :) ?

OpenStudy (kaylala):

okay thanks. so it is true

OpenStudy (***[isuru]***):

u r welcome... and yes the statement is true!